# LLaVA Needs More Knowledge: Retrieval Augmented Natural Language Generation with Knowledge Graph for Explaining Thoracic Pathologies

### v1

@@ 1 Introduction
In recent years, natural language processing (NLP) has witnessed the development of numerous models trained on vast amounts of general domain data [23, 3, 27, 10, 22] While these models exhibit remarkable capabilities across various tasks, they often lack the specialized knowledge required for domain-specific applications, such as generating Natural Language Explanations (NLEs) for thoracic pathologies. This limitation is particularly pronounced in the medical domain, where accurate and contextually relevant explanations are crucial for diagnostic decision-making.
To bridge this gap, including Pre-Training, Fine-Tuning, and Retrieval-Augmented Generation (RAG) methods has been explored. One popular approach among these strategies is Pre-Training large models on medical data. However, this strategy necessitates extensive data and substantial computational resources [26, 16]. In the medical domain, publicly available datasets are often limited, and concerns about data privacy, authenticity, and potential data leakage persist. Even when models are pre-trained on medical data, they frequently struggle with task-specific performance and may suffer from reduced factual accuracy. Moreover, pre-training these models is both expensive and time-consuming, and subsequent fine-tuning is often required to adapt them to specific downstream tasks.
Fine-tuning is another common strategy, where general domain models are adapted directly to medical tasks by training on specialized datasets. While this approach can be effective, it is hampered by the scarcity of high-quality medical datasets and the need to protect patient privacy and data security. Additionally, fine-tuned models can be prone to hallucination, generating explanations that lack factual correctness. Recent advancements in parameter-efficient fine-tuning, such as training low-rank adapters [6], have aimed to reduce computational costs while maintaining model performance, yet challenges remain, particularly in maintaining the model’s generalization across diverse tasks.
The third strategy, RAG [15], has gained traction as a method for enhancing general domain models with domain-specific knowledge dynamically. In this approach, models are fine-tuned on task-specific data while being supplemented with relevant information retrieved from a datastore. RAG methods have shown promising results in maintaining factual accuracy and reducing hallucination risks. However, the effectiveness of RAG is highly dependent on the quality of the retrieval mechanism. Also, in the medical domain, concerns about data privacy are amplified, as retrieved information might still be traceable to individual patients, even after de-identification, thus posing a risk of data leakage.
@@ 2 Related Work
2.1 Natural Language Explanation
@@
NLEs provide textual interpretations of deep learning model predictions, aiming to offer accessible and comprehensible insights for users, particularly in complex domains like medical diagnostics. Hendricks et al. was the first to introduce the NLE task. This task was later extended to encompass the vision-language domain [13, 18, 19, 21, 30]. Kayser et al. introduced the MIMIC-NLE dataset, derived from the MIMIC-CXR dataset [11], to advance interpretability and accessibility in the context of chest X-ray analysis. This dataset is currently the only publicly available resource specifically designed for generating NLEs related to chest X-rays.
@@
In their work, Kayser et al. also introduced benchmark methods for generating explanations, such as DPT (DenseNet-121 [7] combined with GPT-2 [23]) and RATCHET [5]. While GPT-2 has demonstrated effective performance in general domains [13], it shows limitations when applied to medical NLE generation due to its reliance on commonsense knowledge, which is often insufficient for specialized medical contexts. The DPT model, in particular, struggled with generating accurate explanations for chest X-ray images due to its dependency on non-specialized knowledge sources. The MIMIC-NLE dataset provides explanations for predicted pathologies, making it a comprehensive resource for evaluating the quality of NLEs in medical imaging. Our approach demonstrates a substantial improvement over previous methods, underscoring the value of integrating domain-specific knowledge through KG augmentation in the generation of NLEs. Also, Rio-Torto et al. have investigated parameter-efficient training techniques for the NLE task.
2.2 Vision-Language Models
@@
The emergence of advanced Large Language Models (LLMs) like LLaMA [27] and GPT-4 [1] has showcased significant improvements in text generation capabilities. Building on these developments, researchers have increasingly focused on extending these models to handle multimodal inputs, such as visual data. Despite these efforts, fully integrating visual and textual modalities remains a challenging endeavor, particularly in areas such as understanding spatial relationships, mathematical reasoning, and counting. Bordes et al. categorize VLMs into four primary categories, contrastive training, masking, pre-trained backbones, and generative vision-language models.
@@
VLMs in the pre-trained backbone category often utilize open-source LLMs, such as LLaMA/Viccuna, to learn mappings between a pre-trained image encoder and the LLM. This approach is computationally efficient, as it avoids training both text and image encoders from scratch, instead focusing on aligning the representations generated by these pre-trained models [16, 17, 32].
@@
Our work falls within the category of VLMs based on pre-trained backbones. Specifically, we employ an LLaVA [16] model which is based on pre-trained LLM alongside a pre-trained vision model, connected via a Projector that learns to map the visual information to the language model space. This approach leverages the strengths of pre-trained vision and language models.
@@
Visual Instruction Tuning has been effectively employed in models like LLAVA [16], where a pre-trained LLM is trained on visual inputs provided by a vision model using carefully curated instruction datasets. This process typically involves mapping image features from pre-trained models like CLIP into the LLM’s embedding space. While more sophisticated projection methods, such as lightweight transformers in CLIPCap [20] and the Bert-based Q-Former in BLIP-2 [17], have been explored, the linear layer approach remains a popular choice due to its simplicity and efficiency.
2.3 Knowledge Graph
@@
A KG represents relationships between a set of entities, offering a structured approach to capturing and utilizing interconnected information. In the radiology domain, the RadGraph [9] method introduces a pioneering approach to constructing KGs from medical reports, providing a systematic means of extracting and representing clinically relevant information. RadGraph not only defines the methodology for creating these graphs but also supplies a publicly available dataset, which significantly aids research and development in this area.
2.4 Retrieval Augmented Generation
@@
RAG [15] represents a significant advancement in enhancing language models beyond the capabilities achieved through traditional supervised fine-tuning. In RAG, external knowledge is dynamically retrieved and integrated into the language model, thereby augmenting its generative capabilities with new information that the model has not been explicitly trained on. This approach is particularly valuable when addressing dynamic or domain-specific knowledge, such as medical information, where up-to-date and specialized content is crucial. However, as identified by [31], the use of RAG in sensitive domains like healthcare raises substantial security concerns. Specifically, there is a risk that malicious actors could exploit RAG systems by prompting the model to retrieve and expose sensitive information from the retrieval datastore, potentially leading to privacy breaches. This vulnerability underscores the necessity for more secure implementations of RAG, particularly in the context of handling confidential medical data.
@@ 3 Methodology
3.1 Pathology Classification
For the task of pathology classification, which entails predicting the presence of pathologies in X-ray images along with their associated certainty levels, we adopted a methodology based on the approach outlined by [14]. This method involves the prediction of 10 distinct pathologies, each categorized into three certainty levels—negative, uncertain, or positive—using the UMultiClass strategy introduced by [8]. The UMultiClass strategy is specifically designed to address the multi-label nature of pathology classification by effectively categorizing each pathology according to its likelihood of presence. In our work, we used the method employed by [14]. Specifically, we process the visual features extracted from the medical vision model through a multi-layer perceptron (MLP). This modification is intended to enhance the model’s capability to interpret and classify the visual features of the X-ray images, thereby improving the accuracy of both the pathology predictions and their corresponding certainty levels.
3.2 Knowledge Graph Retrieval
@@
To address the risks highlighted in Section 2, we propose a KG-based RAG approach as a more secure alternative. Unlike traditional RAG systems, which may retrieve information that could be traced back to specific patient data, our approach employs a KG composed of general medical terms, entities, and their interrelationships. This structured representation of knowledge abstracts away direct patient-specific details, significantly reducing the risk of inadvertently exposing sensitive information.
@@
In our framework, we augment the model with knowledge instances retrieved from a carefully constructed datastore, conditioning the model to generate responses based on this securely augmented information. This method not only enhances the model’s performance by providing relevant and contextual knowledge but also ensures the protection of sensitive medical data. To enable effective knowledge retrieval, we constructed a datastore comprising KG triplets derived from the MIMIC-CXR training set [11] using RadGraph model [9]. We specifically utilized only the triplets with a ’suggestive_of’ relationship, as these triplets are more directly relevant to explaining the presence of pathologies. These triplet embeddings were generated using a medical CLIP model and are exclusively stored in the datastore, deliberately excluding any image features. This design facilitates a cross-modal retrieval process, where images can be used to query and retrieve relevant text-based knowledge.
For each query image, visual features are extracted using the medical CLIP model, which is designed to map both visual and triplet features into a unified feature space. This unified space allows for the effective retrieval of relevant triplets based on the visual features extracted from the datastore. The retrieval process is conducted by calculating the cosine similarity between the visual features of the query image and the stored triplet embeddings. This enables the selection of the top-k most similar triplets from the KG datastore, ensuring that the retrieved knowledge is closely aligned with the visual content of the query image.
3.3 Vision Models
In our proposed framework, we employ two distinct vision models tailored for different components of the system. First, we utilize the MedCLIP model [29], which serves dual purposes. It is integrated with the DPT [14] framework for NLE generation and is also instrumental in constructing the KGR datastore. MedCLIP’s robust capabilities enable effective retrieval of relevant information based on image features, ensuring that the retrieved knowledge aligns with the visual content of medical images.
For the KGR process within our LLaVA-based framework, we again utilize MedCLIP to perform the retrieval itself based on the image features. However, for the extraction of visual features and their subsequent projection into the language embedding space, we employ the ViT-L/14 [24] CLIP model as the vision encoder. The ViT-L CLIP model extracts visual features from input images, which are then projected into the language embedding space through a trainable projection matrix. Specifically, a simple linear layer is applied to convert the visual features into language embedding tokens, facilitating seamless integration with the language model. This dual-model approach allows us to leverage the specific strengths of both MedCLIP and ViT-L/14 within the framework, optimizing the generation of NLEs and ensuring accurate and contextually relevant retrieval of knowledge.
3.4 Language Models
For the decoding mechanism, we integrated language models such as GPT-2 [23] and LLaMA/Viccuna [27, 22], known for their effectiveness in natural language generation (NLG) tasks. To address the dimensional incompatibility between the CLIP encoder outputs and the language model input requirements, we employed a projector.
3.5 Projector
The GPT-2 based model uses a transformer-based projector, while the LLaMA/Viccuna-based model uses an MLP-based projector. The projector aligns the image embeddings with the features required by the Language Model (LM). We feed the LM with pathologies and their certainty levels (uncertain, positive), along with the retrieved knowledge. These elements are integrated into a structured prompt template. This prompt is then fed into the decoder, which generates NLEs conditioned on the image features, pathologies, and retrieved knowledge. This process ensures the generation of accurate and relevant NLEs tailored to each specific medical case, combining visual data and textual information effectively.
@@ 4 Experiment
4.1 Dataset
For our study, we utilized the MIMIC-NLE dataset [14], which is derived from MIMIC-CXR chest X-ray dataset [11]. This is currently the only publically available dataset for chest X-ray NLEs. The MIMIC-NLE dataset includes diagnoses, evidence labels, and corresponding NLEs for those diagnoses. For a detailed description of the dataset, please refer to the comprehensive overview provided in [14]. The dataset consists of 38,003 NLEs and is divided into training, validation, and testing subsets, containing 37,016, 273, and 714 entries, respectively.
@@ 4.1 Dataset
For our study, we constructed an instruction-format dataset based on the LLaVA [16] framework, with specific modifications to the prompts. These modifications were designed to tailor the questions to elicit detailed explanations from the model regarding the reasoning behind the occurrence of specific pathologies observed in the images. The following questions were incorporated into our prompts to guide the model in generating explanations: Which signs show that the patient has pathologies?, Explain why these pathologies are present in the image?, What evidence in the image indicates pathologies?, How can you tell that the patient has pathologies from the image?, What features suggest the presence of pathologies in this image?. These tailored prompts were crucial in enhancing the model’s ability to provide reasoning for the presence of specific pathologies based on the visual evidence present in the images.
4.2 Training
@@
The KGR process is powered by the MedCLIP model, which projects the input image into a shared latent space. This projected representation is then used to retrieve corresponding triplets from the KG using the FAISS library [12], which employs a k-nearest neighbors algorithm to identify the most relevant triplets. These retrieved triplets serve as supplementary information, enriching the input to the language model and improving the model’s ability to generate accurate and contextually relevant explanations.
@@
RATCHET [5] 66.4 4.7 14.1 22.2 37.9
TieNet [28] 64.6 3.5 12.4 19.4 33.9
DPT [14] 62.5 2.4 11.3 15.4 17.4
LoRA AE [25] 63.9 4.0 15.3 20.6 24.4
Prompt AE [25] 61.3 3.7 14.4 19.7 23.4
Prefix AE [25] 65.2 3.7 14.7 19.7 21.5
LLaMA-Adapt AE [25] 63.9 4.3 14.6 21.4 29.7
+ multi-modal [25] 64.9 3.0 14.1 18.6 18.4
+ MSE loss [25] 62.3 2.0 10.0 14.2 14.2
@@
Table 1: Comparison of our KG-LLaVA framework between other baselines on the MIMIC-NLE test set, focusing on NLG metrics. The metrics include Area Under the Curve (AUC), BLEU-4 (B4), METEOR (MET.), ROUGE-L (R.L.), and CIDEr scores.
Training Setup. We trained the LLaVA [16] model with the integrated KGR module on the MIMIC-NLE dataset. The training process involved fine-tuning the model to maximize the likelihood of generating high-quality NLEs using the additional knowledge provided by the KG.
The training was conducted with LoRA [6] (Low-Rank Adaptation) for training efficiency. We used the pre-trained LLaVA [16] model as the baseline, with a learning rate of 2e-4, and employed the cosine learning rate scheduler. The training was performed over 5 epochs with a batch size of 8 per device and a gradient accumulation of 2 steps. The vision encoder used was CLIP ViT-L [24], and the multimodal projector was an MLP with two layers using the GELU activation function.
In addition to our KG-LLaVA framework, we also applied our KGR module to Med-XPT architecture, which utilizes MedCLIP as the vision encoder, a transformer-based projector, and GPT-2 as the language model. We followed a similar structured training approach. We again used the FAISS [12] library for retrieval based on image features. The training script included several key parameters: the visual encoder (medclip-vit) was used to retrieve captions stored in a JSON file, with the retrieval results incorporated into the input prompts. The attention mechanism within the model was tuned using a cross-attention size of 7, and we enabled the training of the decoder along with the attention mechanism. The training process spanned 15 epochs with a learning rate of 1e-4 and a batch size of 8, utilizing gradient accumulation with one step. The retrieved captions and templates were dynamically integrated into the training pipeline, which facilitated the model’s ability to generate accurate NLEs based on the retrieved knowledge.

### v2

@@ 1 Introduction
In recent years, natural language processing (NLP) has witnessed the development of numerous models trained on vast amounts of general domain data (Radford et al. 2019; Brown et al. 2020; Touvron et al. 2023; Jiang et al. 2023; Peng et al. 2023) While these models exhibit remarkable capabilities across various tasks, they often lack the specialized knowledge required for domain-specific applications, such as generating Natural Language Explanations (NLEs) for thoracic pathologies. This limitation is particularly pronounced in the medical domain, where accurate and contextually relevant explanations are crucial for diagnostic decision-making.
To bridge this gap, including Pre-Training, Fine-Tuning, and Retrieval-Augmented Generation (RAG) methods has been explored. One popular approach among these strategies is Pre-Training large models on medical data. However, this strategy necessitates extensive data and substantial computational resources (Saab et al. 2024; Li et al. 2024). In the medical domain, publicly available datasets are often limited, and concerns about data privacy, authenticity, and potential data leakage persist. Even when models are pre-trained on medical data, they frequently struggle with task-specific performance and may suffer from reduced factual accuracy. Moreover, pre-training these models is both expensive and time-consuming, and subsequent fine-tuning is often required to adapt them to specific downstream tasks.
Fine-tuning is another common strategy, where general domain models are adapted directly to medical tasks by training on specialized datasets. While this approach can be effective, it is hampered by the scarcity of high-quality medical datasets and the need to protect patient privacy and data security. Additionally, fine-tuned models can be prone to hallucination, generating explanations that lack factual correctness. Recent advancements in parameter-efficient fine-tuning, such as training low-rank adapters (Hu et al. 2022), have aimed to reduce computational costs while maintaining model performance, yet challenges remain, particularly in maintaining the model’s generalization across diverse tasks.
The third strategy, RAG (Lewis et al. 2020), has gained traction as a method for enhancing general domain models with domain-specific knowledge dynamically. In this approach, models are fine-tuned on task-specific data while being supplemented with relevant information retrieved from a datastore. RAG methods have shown promising results in maintaining factual accuracy and reducing hallucination risks. However, the effectiveness of RAG is highly dependent on the quality of the retrieval mechanism. Also, in the medical domain, concerns about data privacy are amplified, as retrieved information might still be traceable to individual patients, even after de-identification, thus posing a risk of data leakage.
@@ 2 Related Work
Natural Language Explanation.
@@
NLEs provide textual interpretations of deep learning model predictions, aiming to offer accessible and comprehensible insights for users, particularly in complex domains like medical diagnostics. Hendricks et al. was the first to introduce the NLE task. This task was later extended to encompass the vision-language domain (Kayser et al. 2021; Li et al. 2018; Marasović et al. 2020; Park et al. 2018; Wu and Mooney 2019). Kayser et al. introduced the MIMIC-NLE dataset, derived from the MIMIC-CXR dataset (Johnson et al. 2019), to advance interpretability and accessibility in the context of chest X-ray analysis. This dataset is currently the only publicly available resource specifically designed for generating NLEs related to chest X-rays.
@@
In their work, Kayser et al. also introduced benchmark methods for generating explanations, such as DPT (DenseNet-121 (Huang et al. 2017) combined with GPT-2 (Radford et al. 2019)) and RATCHET (Hou et al. 2021). While GPT-2 has demonstrated effective performance in general domains (Kayser et al. 2021), it shows limitations when applied to medical NLE generation due to its reliance on commonsense knowledge, which is often insufficient for specialized medical contexts. The DPT model, in particular, struggled with generating accurate explanations for chest X-ray images due to its dependency on non-specialized knowledge sources. The MIMIC-NLE dataset provides explanations for predicted pathologies, making it a comprehensive resource for evaluating the quality of NLEs in medical imaging. Our approach demonstrates a substantial improvement over previous methods, underscoring the value of integrating domain-specific knowledge through KG augmentation in the generation of NLEs. Also, Rio-Torto, Cardoso, and Teixeira have investigated parameter-efficient training techniques for the NLE task.
Vision-Language Models.
@@
The emergence of advanced Large Language Models (LLMs) like LLaMA (Touvron et al. 2023) and GPT-4 (Achiam et al. 2023) has showcased significant improvements in text generation capabilities. Building on these developments, researchers have increasingly focused on extending these models to handle multimodal inputs, such as visual data. Despite these efforts, fully integrating visual and textual modalities remains a challenging endeavor, particularly in areas such as understanding spatial relationships, mathematical reasoning, and counting. Bordes et al. categorize VLMs into four primary categories, contrastive training, masking, pre-trained backbones, and generative vision-language models.
@@
VLMs in the pre-trained backbone category often utilize open-source LLMs, such as LLaMA/Viccuna, to learn mappings between a pre-trained image encoder and the LLM. This approach is computationally efficient, as it avoids training both text and image encoders from scratch, instead focusing on aligning the representations generated by these pre-trained models (Li et al. 2024, 2023; Zhu et al. 2024).
@@
Our work falls within the category of VLMs based on pre-trained backbones. Specifically, we employ an LLaVA (Li et al. 2024) model which is based on pre-trained LLM alongside a pre-trained vision model, connected via a Projector that learns to map the visual information to the language model space. This approach leverages the strengths of pre-trained vision and language models.
@@
Visual Instruction Tuning has been effectively employed in models like LLAVA (Li et al. 2024), where a pre-trained LLM is trained on visual inputs provided by a vision model using carefully curated instruction datasets. This process typically involves mapping image features from pre-trained models like CLIP into the LLM’s embedding space. While more sophisticated projection methods, such as lightweight transformers in CLIPCap (Mokady, Hertz, and Bermano 2021) and the Bert-based Q-Former in BLIP-2 (Li et al. 2023), have been explored, the linear layer approach remains a popular choice due to its simplicity and efficiency.
Knowledge Graph.
@@
A KG represents relationships between a set of entities, offering a structured approach to capturing and utilizing interconnected information. In addition to KG-TREAT (Liu, Wu, and Zhang 2024) and MKG-FENN (Wu et al. 2024), recent works such as ImgFact (Liu et al. 2024) and VCTP (Chen et al. 2024) have pushed the boundaries of KG integration. ImgFact introduces a large-scale multimodal KG with triplet fact grounding to improve visual reasoning tasks, while VCTP applies visual chain-of-thought prompting to facilitate knowledge-based visual reasoning. In the radiology domain, the RadGraph (Jain et al. 2021) method introduces a pioneering approach to constructing KGs from medical reports, providing a systematic means of extracting and representing clinically relevant information. RadGraph not only defines the methodology for creating these graphs but also supplies a publicly available dataset, which significantly aids research and development in this area.
Unlike these methods, our privacy-preserving KG-RAG framework focuses on medical NLE tasks. By leveraging clinically relevant relationships (e.g., ‘suggestive_of’) and ensuring data privacy through de-identified triplets for NLE generation.
Retrieval Augmented Generation.
@@
RAG (Lewis et al. 2020) represents a significant advancement in enhancing language models beyond the capabilities achieved through traditional supervised fine-tuning. In RAG, external knowledge is dynamically retrieved and integrated into the language model, thereby augmenting its generative capabilities with new information that the model has not been explicitly trained on. This approach is particularly valuable when addressing dynamic or domain-specific knowledge, such as medical information, where up-to-date and specialized content is crucial. However, as identified by (Zeng et al. 2024), the use of RAG in sensitive domains like healthcare raises substantial security concerns. Specifically, there is a risk that malicious actors could exploit RAG systems by prompting the model to retrieve and expose sensitive information from the retrieval datastore, potentially leading to privacy breaches. This vulnerability underscores the necessity for more secure implementations of RAG, particularly in the context of handling confidential medical data.
@@ 3 Methodology
Pathology Classification
For pathology classification, the input X-ray image $X\in\mathbb{R}^{H\times W\times C}$ is processed by a visual encoder, such as MedCLIP or ViT-L/14, to generate visual feature embeddings:
$Z_{v}=V(X),$ (1)
where $V(X)$ represents the visual encoder that maps $X$ to a high-dimensional feature vector $Z_{v}\in\mathbb{R}^{d_{v}}$.
The extracted visual features $Z_{v}$ are then passed through a multi-layer perceptron (MLP) to classify pathologies:
$\hat{y}=\text{MLP}(Z_{v}),$ (2)
where $\hat{y}\in\mathbb{R}^{p}$ denotes the predicted probabilities for $p$ pathology classes.
In addition to predicting the associated certainty levels along with the identified pathologies in X-ray images, we adopted a methodology based on the approach outlined by (Kayser et al. 2022). This method involves the prediction of 10 distinct pathologies, each categorized into three certainty levels—negative, uncertain, or positive—using the UMultiClass strategy introduced by (Irvin et al. 2019).
[…]
$c_{i}=f(\hat{y}_{i})=\begin{cases}\text{negative}&\text{if }\hat{y}_{i}<\theta_{\text{neg}},\\ \text{uncertain}&\text{if }\theta_{\text{neg}}\leq\hat{y}_{i}<\theta_{\text{pos}},\\ \text{positive}&\text{if }\hat{y}_{i}\geq\theta_{\text{pos}},\end{cases}$ (3)
where $\theta_{\text{neg}}$ and $\theta_{\text{pos}}$ are the thresholds for classifying certainty levels.
We process the visual features $Z_{v}$ extracted from the medical vision model through the MLP to enhance the model’s capability to interpret and classify the visual features of the X-ray images. This process improves the accuracy of both the pathology predictions $\hat{y}$ and their corresponding certainty levels $c_{i}$.
Knowledge Graph Retrieval
The Knowledge Graph (KG) is represented as a collection of triplets $(e_{i},r_{ij},e_{j})$, where $e_{i}$ and $e_{j}$ are medical entities, and $r_{ij}$ is the relationship between them extracted from clinical reports:
$\text{KG}=\{(e_{i},r_{ij},e_{j})\mid e_{i},e_{j}\in\mathcal{E},r_{ij}\in\mathcal{R}\}.$ (4)
Each triplet is embedded into a high-dimensional feature space using the MedCLIP text encoder $f_{KG}$:
$Z_{k}=f_{KG}(e_{i},r_{ij},e_{j}),$ (5)
where $Z_{k}\in\mathbb{R}^{d_{k}}$ represents the embedding of the triplet $(e_{i},r_{ij},e_{j})$.
To retrieve the most relevant triplets for a given query image $X$, the image is first processed by the visual encoder to obtain the image embedding (refer to Equation 1).
The retrieval process is conducted by computing the cosine similarity between the image embedding $Z_{v}$ and each triplet embedding $Z_{k}$ in the datastore:
$\text{sim}(Z_{v},Z_{k})=\frac{Z_{v}\cdot Z_{k}}{\|Z_{v}\|\|Z_{k}\|}.$ (6)
The top-$k$ most relevant triplets $T_{\text{retrieved}}$ are then selected based on the highest similarity scores:
$T_{\text{retrieved}}=\arg\max_{Z_{k}\in\text{KG}}\,\text{sim}(Z_{v},Z_{k}).$ (7)
@@
To address the privacy risks highlighted in Section 2, we propose a KG-based RAG approach as a secure alternative. Unlike traditional RAG systems, which may retrieve information traceable to specific patients, our approach employs a KG composed of general medical terms, entities, and relationships. This structured representation of knowledge graph abstracts away direct patient-specific details, significantly reducing the risk of inadvertently exposing sensitive information.
@@
In our framework, we augment the model with knowledge instances retrieved from a constructed datastore. This method enhances the model’s performance by providing relevant contextual knowledge while protecting sensitive medical data. To enable effective knowledge retrieval, we constructed a datastore comprising KG triplets derived exclusively from the MIMIC-CXR training set (Johnson et al. 2019) using the RadGraph model (Jain et al. 2021). These triplets are aligned with the MIMIC-NLE (Kayser et al. 2022) training set, ensuring no overlap with the dev and test sets, thereby preventing data leakage. We focus specifically on triplets with the ‘suggestive_of’ relationship (e.g., “opacity suggestive_of pneumonia”) as these triplets are more directly relevant to explaining the presence of pathologies. These triplet embeddings were generated using a medical CLIP model and are exclusively stored in the datastore, deliberately excluding any image features. This design facilitates a cross-modal retrieval process, where images can be used to query and retrieve relevant KG triplets.
KG-RAG retrieves contextually relevant knowledge by focusing on semantically similar information rather than exact matches. Even when minor deviations occur, the Vision-Language Model (VLM) leverages visual cues to maintain accurate pathology identification.
@@ 4 Experiment
Dataset
For our study, we utilized the MIMIC-NLE dataset (Kayser et al. 2022), which is derived from MIMIC-CXR chest X-ray dataset (Johnson et al. 2019). This is currently the only publically available dataset for chest X-ray NLEs. The MIMIC-NLE dataset includes diagnoses, evidence labels, and corresponding NLEs for those diagnoses. For a detailed description of the dataset, please refer to the comprehensive overview provided in (Kayser et al. 2022). The dataset consists of 38,003 NLEs and is divided into training, validation, and testing subsets, containing 37,016, 273, and 714 entries, respectively.
@@
For a comprehensive description of the instruction-format dataset and the tailored prompts for pathology explanations, please refer to Appendix A in the supplementary material.
Implementation Details
Projector. The GPT-2 based model uses a transformer-based projector, while the LLaMA/Vicuna-based model uses an MLP-based projector. The projector aligns the image embeddings with the features required by the Language Model (LM). Specifically, the visual features $Z_{v}$ are projected into the language model’s embedding space using a trainable linear layer:
$H_{v}=W_{p}\cdot Z_{v}+b_{p},$ (8)
where $W_{p}\in\mathbb{R}^{d_{l}\times d_{v}}$ is the weight matrix, $b_{p}\in\mathbb{R}^{d_{l}}$ is the bias term, and $H_{v}\in\mathbb{R}^{d_{l}}$ represents the projected visual features aligned with the language model’s embedding space.
Vision Models. In our proposed framework, we employ two distinct vision models tailored for different components of the system. First, we utilize the MedCLIP model (Wang et al. 2022), which serves dual purposes. It is integrated with the DPT (Kayser et al. 2022) framework for NLE generation and is also instrumental in constructing the KGR datastore. MedCLIP’s robust capabilities enable effective retrieval of relevant information based on image features, ensuring that the retrieved knowledge aligns with the visual content of medical images.
For the KGR process within our LLaVA-based framework, we again utilize MedCLIP to perform the retrieval itself based on the image features. However, for the extraction of visual features and their subsequent projection into the language embedding space, we employ the ViT-L/14 (Radford et al. 2021) CLIP model as the vision encoder. The ViT-L CLIP model extracts visual features from input images (refer to Equation 1).
These visual features are then projected into the language model’s embedding space through a trainable projection matrix (refer to Equation 8). This simple linear layer facilitates seamless integration with the language model. This dual-model approach allows us to leverage the specific strengths of both MedCLIP and ViT-L/14 within the framework, optimizing the generation of NLEs and ensuring accurate and contextually relevant retrieval of knowledge.
Language Models. For the decoding mechanism, we integrated language models such as GPT-2 (Radford et al. 2019) and LLaMA/Viccuna (Touvron et al. 2023; Peng et al. 2023), known for their effectiveness in natural language generation (NLG) tasks. The language model $LM$ generates a probability distribution over target tokens $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ conditioned on the input sequence $S$, which includes projected visual features, pathology labels, and retrieved knowledge triplets:
$P(Y\mid S)=\prod_{i=1}^{n}P(y_{i}\mid S,y_{1},y_{2},\ldots,y_{i-1}).$ (9)
This structured prompt ensures that the generated NLEs are conditioned on relevant image features, pathology information, and retrieved knowledge. This process effectively combines visual data and textual information to produce accurate and contextually relevant NLEs.
Training
@@
RATCHET (Hou et al. 2021) 66.4 4.7 14.1 22.2 37.9
TieNet (Wang et al. 2018) 64.6 3.5 12.4 19.4 33.9
DPT (Kayser et al. 2022) 62.5 2.4 11.3 15.4 17.4
LoRA AE (Rio-Torto, Cardoso, and Teixeira 2024) 63.9 4.0 15.3 20.6 24.4
Prompt AE (Rio-Torto, Cardoso, and Teixeira 2024) 61.3 3.7 14.4 19.7 23.4
Prefix AE (Rio-Torto, Cardoso, and Teixeira 2024) 65.2 3.7 14.7 19.7 21.5
LLaMA-Adapt AE (Rio-Torto, Cardoso, and Teixeira 2024) 63.9 4.3 14.6 21.4 29.7
+ multi-modal (Rio-Torto, Cardoso, and Teixeira 2024) 64.9 3.0 14.1 18.6 18.4
+ MSE loss (Rio-Torto, Cardoso, and Teixeira 2024) 62.3 2.0 10.0 14.2 14.2
@@
Table 1: Comparison of our KG-LLaVA with other baselines on the MIMIC-NLE test set, focusing on NLG metrics. The metrics include Area Under the Curve (AUC), BLEU-4 (B4), METEOR (MET.), ROUGE-L (R.L.), and CIDEr scores.
@@
Refer to Appendix B in the supplementary material for detailed implementation and training configurations of KG-LLaVA, Med-XPT, and Bio-LLaVA.
Evaluation Metrics
In line with (Kayser et al. 2022), we evaluate NLEs only for correctly predicted labels. Prior research has indicated that standard automated natural language generation (NLG) metrics often fall short in assessing NLE quality due to the variability in expressing similar meanings using different syntactic structures and semantic interpretations (Kayser et al. 2021). For our evaluation, we report the widely used NLG metrics: BLEU, ROUGE, CIDEr, and METEOR (Kayser et al. 2021).
Among these, CIDEr is particularly relevant as it measures the n-gram overlap with ground-truth explanations, placing emphasis on contextual relevance. By retrieving pathology-specific triplets through KG-RAG, our approach increases the likelihood of incorporating relevant medical terms into the generated NLEs. This contributes to the higher CIDEr scores observed in our results, highlighting the effectiveness of KG-RAG in enhancing the contextual accuracy of explanations.